# MixFlow: Mixture-Conditioned Flow Matching for Out-of-Distribution Generalization

### v1

@@ 5.2 Using MixFlow to predict single-cell responses for unseen interventions
Single-cell perturbation screenings are an invaluable tool for new treatment discovery; they consist in testing the effects of genetic or chemical perturbations on cells in vitro. Each cell is exposed to a specific condition, and its gene expression is measured to assess whether the perturbation is associated with a given phenotypic change. We evaluate whether MixFlow can predict cellular responses to perturbations never observed before.
@@
We report Wasserstein-1, Wasserstein-2, MMD, and Energy distances (ED). These metrics assess whether predicted distributions accurately capture the global geometry of observed cell states, beyond low-order statistics. On all tested single-cell perturbation datasets, MixFlow achieves substantially stronger distributional alignment than Vanilla CFM (Table 3), showing how MixFlow better generalises when generating previously unseen conditions, across both genetic and chemical perturbation modalities.
Table 2: Single-cell perturbation datasets used to evaluate MixFlow.
@@
Table 3: Cell-type perturbation benchmarks comparing MixFlow and CFM. Mean $\pm$ std over four runs. Lower is better.
@@
MixFlow 0.095 $\pm$ 0.003 52.868 $\pm$ 0.215 1404.9 $\pm$ 11.5 27.3 $\pm$ 0.4
5.3 MixFlow predicts unseen intervention responses in image-based assays
To extend beyond transcriptomic data, we evaluate MixFlow in an image-based drug discovery setting. High-content imaging provides rich phenotypic readouts, but exhaustive compound testing is infeasible. We therefore train on a subset of compounds and predict cell morphologies for unseen compounds using only their chemical representations at test time.
We benchmark on BBBC021 Caie et al. (2010) and RxRx1 Sypetkowski et al. (2023), comparing MixFlow against a conditional flow matching baseline (CFM). Table 4 shows that MixFlow achieves consistently lower distance metrics (MMD, W1, W2, ED), indicating more faithful recovery of unseen phenotypic responses. We split BBBC021 following prior work (Palma et al., 2025): training uses 25 compounds and testing 9. For RxRx1, we train on 853 siRNAs and test on 213.
A key design choice is to generate instances in feature space rather than pixel space. Images are encoded with DINOv2 (Oquab et al., 2024) into $768$-D embeddings, which aligns with semantically meaningful morphology-level distances (Oquab et al., 2024; Zhang et al., 2018). This strategy improves efficiency, imposes a semantic inductive bias, and enables geometry-aware evaluation. For fairness, CFM outputs are also scored in the same DINOv2 feature space. Results are provided in Table 4. These substantial gains indicate that MixFlow more faithfully recovers unseen compound phenotypes, highlighting its potential to reduce experimental burden through reliable in silico prediction.
Table 4: BBBC021 and RxRx1 benchmarks comparing MixFlow and CFM on phenotypic perturbations. Lower is better; values are the mean over three runs.
@@ 6 Ablations
We conduct ablation studies to understand the key design choices in MixFlow. All experiments are performed on both transcriptomic (iAstrocytes) and morphological (BBBC021) datasets to ensure findings generalize across modalities.
@@
A central design choice in MixFlow is the number of modes $I$ in the Gaussian mixture base distribution. We vary $I$ (the number of modes) and measure the energy distance between generated and ground-truth distributions on held-out perturbations. As shown in Figure 3(a), increasing the number of modes consistently improves performance up to a point: energy distance decreases as $I$ grows, reflecting the increased expressiveness of the base distribution. However, beyond a dataset-dependent threshold, performance plateaus or slightly degrades. We attribute this to optimization difficulty: with too many modes, the model must learn to allocate probability mass across components that may not be necessary, increasing the effective number of parameters without proportional benefit. Notably, the optimal $I$ differs between datasets (transcriptomic data benefits from more modes than morphological data) suggesting that the appropriate complexity of the base distribution should reflect the heterogeneity of the target populations. These findings align with our theoretical analysis in Section 4: increasing $I$ reduces the degrees of freedom in the transport problem by $D$ per mode, improving identifiability (Statement 1, §2).
@@
A known challenge for flow-based models is scaling to high-dimensional spaces, where the fixed Gaussian base distribution $\mathcal{N}(0,I)$ becomes increasingly mismatched to structured target distributions. We investigate this by training both MixFlow and CFM on PCA-reduced representations with increasing numbers of components. Figure 3(b) shows that CFM’s performance degrades substantially as dimensionality increases: the energy distance grows by over an order of magnitude between 50 and 500 PCA components. In contrast, MixFlow maintains stable performance across the full range of dimensionalities tested. This robustness arises because the learned GMM base distribution adapts to the structure of each condition, reducing the effective transport distance even in high dimensions. From the perspective of Section 4, higher $D$ increases the constraint count ($ID+I$), which for fixed $I$ reduces the residual degrees of freedom JIDJ-ID, explaining why MixFlow’s advantage becomes more pronounced in higher dimensions.
These ablations confirm two key properties of MixFlow: (i) the number of mixture components provides a useful complexity knob that should be tuned to dataset heterogeneity, and (ii) conditioning the base distribution yields substantial robustness benefits in high-dimensional settings where standard CFM struggles.
@@
(b) Energy distance versus PCA dimensionality for MixFlow and CFM. MixFlow maintains stable performance while CFM degrades in higher dimensions.
Figure 3: Ablation studies on transcriptomic (Combo-SciPlex) and morphological (BBBC021) datasets. (a) Effect of the number of GMM modes on generation quality. (b) Comparison of MixFlow and CFM robustness across increasing data dimensionality.
@@ 7 Limitations and Future Work
While MixFlow demonstrates strong performance across diverse perturbation modalities, several limitations remain. The framework depends on the availability of high-quality perturbation descriptors such as gene embeddings or chemical fingerprints; incomplete or noisy metadata may hinder predictive accuracy. Our evaluation focuses primarily on single perturbations and small combinations, whereas scaling to higher-order combinations, dose–response landscapes, and time-course dynamics remains an open challenge. Moreover, although we provide theoretical justification for out-of-distribution generalisation, the analysis assumes smooth transport fields that may not fully capture heterogeneous tissues or disease states.
Future work should address these challenges by incorporating richer multimodal descriptors, extending the model to capture interaction effects and temporal responses, and scaling training to very large datasets. In particular, applying MixFlow to resources such as the JUMP- consortium dataset for imaging assays (Chandrasekaran et al., 2024) and large single-cell perturbation compendia (Zhang et al., 2025a) could enable the development of true zero-shot models that leverage perturbation similarity at scale. Finally, extending the framework to translational settings such as patient-derived cells or organoids will be critical for realising its potential impact in therapeutic design.

### v2

@@ 5.2 Using SP-FM to predict single-cell responses for unseen interventions
Single-cell perturbation screenings are an invaluable tool for new treatment discovery; they consist in testing the effects of genetic or chemical perturbations on cells in vitro. Each cell is exposed to a specific condition, and its gene expression is measured to assess whether the perturbation is associated with a given phenotypic change. We evaluate whether SP-FM can predict cellular responses to perturbations never observed before.
@@
We report Wasserstein-1, Wasserstein-2, MMD, and Energy distances (ED). These metrics assess whether predicted distributions accurately capture the global geometry of observed cell states, beyond low-order statistics. On all tested single-cell perturbation datasets, SP-FM achieves substantially stronger distributional alignment than Vanilla CFM (Table 3), showing how SP-FM better generalises when generating previously unseen conditions, across both genetic and chemical perturbation modalities.
Table 2: Single-cell perturbation datasets used to evaluate SP-FM.
@@
Table 3: Cell-type perturbation benchmarks comparing SP-FM and CFM. Mean $\pm$ std over four runs. Lower is better.
@@
SP-FM 0.095 $\pm$ 0.003 52.868 $\pm$ 0.215 1404.9 $\pm$ 11.5 27.3 $\pm$ 0.4
5.3 SP-FM predicts unseen intervention responses in image-based assays
To extend beyond transcriptomic data, we evaluate SP-FM in an image-based drug discovery setting. High-content imaging provides rich phenotypic readouts, but exhaustive compound testing is infeasible. We therefore train on a subset of compounds and predict cell morphologies for unseen compounds using only their chemical representations at test time.
We benchmark on BBBC021 Caie et al. (2010) and RxRx1 Sypetkowski et al. (2023), comparing SP-FM against a conditional flow matching baseline (CFM). Table 4 shows that SP-FM achieves consistently lower distance metrics (MMD, W1, W2, ED), indicating more faithful recovery of unseen phenotypic responses. We split BBBC021 following prior work (Palma et al., 2025): training uses 25 compounds and testing 9. For RxRx1, we train on 853 siRNAs and test on 213.
A key design choice is to generate instances in feature space rather than pixel space. Images are encoded with DINOv2 (Oquab et al., 2024) into $768$-D embeddings, which aligns with semantically meaningful morphology-level distances (Oquab et al., 2024; Zhang et al., 2018). This strategy improves efficiency, imposes a semantic inductive bias, and enables geometry-aware evaluation. For fairness, CFM outputs are also scored in the same DINOv2 feature space. Results are provided in Table 4. These substantial gains indicate that SP-FM more faithfully recovers unseen compound phenotypes, highlighting its potential to reduce experimental burden through reliable in silico prediction.
Table 4: BBBC021 and RxRx1 benchmarks comparing SP-FM and CFM on phenotypic perturbations. Lower is better; values are the mean over three runs.
@@ 6 Ablations
We conduct ablation studies to understand the key design choices in SP-FM. All experiments are performed on both transcriptomic (iAstrocytes) and morphological (BBBC021) datasets to ensure findings generalize across modalities.
@@
A central design choice in SP-FM is the number of modes $I$ in the Gaussian mixture base distribution. We vary $I$ (the number of modes) and measure the energy distance between generated and ground-truth distributions on held-out perturbations. As shown in Figure 3(a), increasing the number of modes consistently improves performance up to a point: energy distance decreases as $I$ grows, reflecting the increased expressiveness of the base distribution. However, beyond a dataset-dependent threshold, performance plateaus or slightly degrades. We attribute this to optimization difficulty: with too many modes, the model must learn to allocate probability mass across components that may not be necessary, increasing the effective number of parameters without proportional benefit. Notably, the optimal $I$ differs between datasets (transcriptomic data benefits from more modes than morphological data) suggesting that the appropriate complexity of the base distribution should reflect the heterogeneity of the target populations. These findings align with our theoretical analysis in Section 4: increasing $I$ reduces the degrees of freedom in the transport problem by $D$ per mode, improving identifiability (Statement 1, §2).
@@
A known challenge for flow-based models is scaling to high-dimensional spaces, where the fixed Gaussian base distribution $\mathcal{N}(0,I)$ becomes increasingly mismatched to structured target distributions. We investigate this by training both SP-FM and CFM on PCA-reduced representations with increasing numbers of components. Figure 3(b) shows that CFM’s performance degrades substantially as dimensionality increases: the energy distance grows by over an order of magnitude between 50 and 500 PCA components. In contrast, SP-FM maintains stable performance across the full range of dimensionalities tested. This robustness arises because the learned GMM base distribution adapts to the structure of each condition, reducing the effective transport distance even in high dimensions. From the perspective of Section 4, higher $D$ increases the constraint count ($ID+I$), which for fixed $I$ reduces the residual degrees of freedom JIDJ-ID, explaining why SP-FM’s advantage becomes more pronounced in higher dimensions.
These ablations confirm two key properties of SP-FM: (i) the number of mixture components provides a useful complexity knob that should be tuned to dataset heterogeneity, and (ii) conditioning the base distribution yields substantial robustness benefits in high-dimensional settings where standard CFM struggles.
@@
(b) Energy distance versus PCA dimensionality for SP-FM and CFM. SP-FM maintains stable performance while CFM degrades in higher dimensions.
Figure 3: Ablation studies on transcriptomic (Combo-SciPlex) and morphological (BBBC021) datasets. (a) Effect of the number of GMM modes on generation quality. (b) Comparison of SP-FM and CFM robustness across increasing data dimensionality.
@@ 7 Limitations and Future Work
While SP-FM demonstrates strong performance across diverse perturbation modalities, several limitations remain. The framework depends on the availability of high-quality perturbation descriptors such as gene embeddings or chemical fingerprints; incomplete or noisy metadata may hinder predictive accuracy. Our evaluation focuses primarily on single perturbations and small combinations, whereas scaling to higher-order combinations, dose–response landscapes, and time-course dynamics remains an open challenge. Moreover, although we provide theoretical justification for out-of-distribution generalisation, the analysis assumes smooth transport fields that may not fully capture heterogeneous tissues or disease states.
Future work should address these challenges by incorporating richer multimodal descriptors, extending the model to capture interaction effects and temporal responses, and scaling training to very large datasets. In particular, applying SP-FM to resources such as the JUMP- consortium dataset for imaging assays (Chandrasekaran et al., 2024) and large single-cell perturbation compendia (Zhang et al., 2025a) could enable the development of true zero-shot models that leverage perturbation similarity at scale. Finally, extending the framework to translational settings such as patient-derived cells or organoids will be critical for realising its potential impact in therapeutic design.
@@ Appendix A Generalisation analysis, details and proofs
Appendix B Letters
B.1 Dataset details
B.2 Synthetic Letters Dataset (Letter–Rotation Benchmark)
Data generation.
We construct a synthetic benchmark by rendering uppercase letters as binary silhouettes and sampling point clouds from the foreground pixels. Each condition corresponds to a (letter, rotation) pair. For a given letter, we generate $R=20$ rotations uniformly spaced in $0,2\pi]$. For each condition, we create multiple replicas by rendering the same silhouette with different RGB colours; these replicas are used only to provide multiple samples from the same conditional distribution and are not included in the conditioning vector.
Point-cloud representation.
Each sample is a point cloud in $\mathbb{R}^{D}$ with $D=5$, consisting of 2D coordinates and RGB channels: $\bm{x}=(x,y,r,g,b).$ We extract foreground-pixel coordinates, normalize them to $1,1]^{2}$, rescale to correct aspect ratio using the glyph’s bounding box, optionally add Gaussian coordinate noise (set to 0 in our experiments), and finally rotate the point cloud by the condition’s angle.
Colour replicas.
For each (letter, rotation) condition, we generate $C$ colour replicas to better approximate the underlying conditional distribution. In our experiments we use $C=100$. To ensure that colour does not dominate the evaluation metrics, we restrict the colour variation to a narrow range by sampling only the red channel $r\sim\mathrm{Unif}(r_{\min},r_{\max})$ with $(r_{\min},r_{\max})=(0.5,0.6)$ and setting $g=b=0$; thus colour provides mild diversity without strongly affecting distance metrics.
Condition descriptor.
The conditioning vector $\bm{y}$ concatenates (i) a one-hot encoding of the letter identity and (ii) a continuous rotation scalar normalized to $[0,1]$:
$\bm{y}=[\mathrm{onehot}(\text{letter});\;\theta/(2\pi)].$
Colour is not included in $\bm{y}$.
Source distribution.
For conditional flow-matching training we sample the source point cloud from an isotropic Gaussian. In the our setting, each source point cloud uses a shared RGB value across all points (sampled uniformly within the same restricted colour range) while the 2D coordinates are sampled from $\mathcal{N}(0,I)$, matching the data format of the target clouds.
Train/validation/test splits (rotation interpolation).
We study rotation interpolation by splitting the rotation indices: training uses even-indexed rotations, while validation and test use odd-indexed rotations. All letters appear in training (but only at even rotations). The validation set uses a designated validation letter (S in the main paper), evaluated on odd rotations. Test sets are formed from unseen test letters (W and Y in the main paper), also evaluated on odd rotations. Additional results with alternative validation/test letters are reported in the Table 5.
Subsampling for OT coupling.
Each rendered silhouette contains many foreground pixels; to make mini-batch OT pairing tractable, we subsample a fixed number of points per cloud. We use $n=100$ points per cloud because the discrete OT coupling has super-cubic complexity in $n$ in our implementation, making larger clouds prohibitively expensive. This subsampling is applied consistently across all methods.
Key hyperparameters.
Unless otherwise stated, we use: batch size $10$, point subsampling $n=100$, number of letters $4$, number of rotations $R=20$, number of colour replicas per condition $C=100$, $(r_{\min},r_{\max})=(0.5,0.6)$, and zero coordinate noise.
B.3 SP-FM implementation
B.4 SP-FM / PertFlow on Colored Letter Point Clouds
Data representation and conditions.
Each sample is a point cloud of $N$ points, where each point carries 2D geometry and RGB color:
$x\;=\;(x,y,r,g,b)\in\mathbb{R}^{5},\qquad\mathbf{X}\in\mathbb{R}^{N\times 5}.$
We condition generation on the letter identity and rotation. The condition vector is
$\mathbf{y}=[\text{one-hot}(\text{letter}),\,\rho]\in\mathbb{R}^{C},$
where $\rho\in[0,1]$ is the normalized rotation angle (mapped to $0,2\pi]$ when decoding for visualization).
Dual-flow architecture (geometry vs. color).
We use a dual-MLP conditional flow architecture: (i) a geometry flow $v_{\text{xy}}$ acting on $(x,y)\in\mathbb{R}^{2}$ per point, and (ii) a color flow $v_{\text{rgb}}$ acting on an RGB vector $(r,g,b)\in\mathbb{R}^{3}$ per cloud. Concretely, we decompose each cloud as
$\mathbf{X}_{\text{xy}}\in\mathbb{R}^{N\times 2},\qquad\mathbf{c}\in\mathbb{R}^{1\times 3},$
where $\mathbf{c}$ denotes the (shared) RGB color of the cloud (in our setup, all points in a cloud share the same color). Both velocity fields are implemented as MLPs that take time $t$ and the concatenation of state and condition:
$v_{\text{xy}}(t,\mathbf{x}_{\text{xy}};\mathbf{y})\;=\;\mathrm{MLP}_{\text{xy}}\!\big(t,\,[\mathbf{x}_{\text{xy}},\mathbf{y}]\big),\qquad v_{\text{rgb}}(t,\mathbf{c};\mathbf{y})\;=\;\mathrm{MLP}_{\text{rgb}}\!\big(t,\,[\mathbf{c},\mathbf{y}]\big).$
Conditional base distribution with random per-condition mode initialization.
To sample the initial state $\mathbf{X}_{0}$ we use a condition-dependent mixture base. In the use_kmeans=True setting, mixture centers are initialized by randomly sampling $K$ training point clouds for each condition $\mathbf{y}$ (i.e., for each letter-rotation pair), rather than running K-means. These sampled clouds serve as fixed centers, while the mixture weights are predicted by a small network $h_{p}(\mathbf{y})$. At sampling time we draw a (soft) mixture over centers via Gumbel–Softmax (temperature $\tau$) and add Gaussian perturbation with variance $\sigma^{2}$ to obtain $\mathbf{X}_{0,\text{xy}}$.
For color, we initialize a single RGB value per cloud by sampling uniformly from a predefined range $a,b]\subset[0,1]$ and broadcasting it to all points:
$\mathbf{c}_{0}\sim\mathcal{U}([a,b]^{3}),\qquad\mathbf{X}_{0,\text{rgb}}=\mathbf{1}_{N}\mathbf{c}_{0},$
and finally concatenate geometry and color to obtain $\mathbf{X}_{0}=[\mathbf{X}_{0,\text{xy}},\mathbf{X}_{0,\text{rgb}}]\in\mathbb{R}^{N\times 5}$.
OT pairing.
Before computing training losses, we align points between the sampled source cloud $\mathbf{X}_{0}$ and the target cloud $\mathbf{X}_{1}$ using point-wise optimal transport (OT) within each sample, which reduces sensitivity to arbitrary point ordering.
Flow matching objective.
Given an OT-aligned pair $(\mathbf{X}_{0},\mathbf{X}_{1})$ under condition $\mathbf{y}$, we sample $t\sim\mathcal{U}[0,1]$ and define
$\mathbf{X}_{t}=(1-t)\mathbf{X}_{0}+t\mathbf{X}_{1},\qquad\mathbf{U}=\mathbf{X}_{1}-\mathbf{X}_{0}.$
We use the standard flow-matching regression form
$\mathcal{L}_{\mathrm{FM}}=\mathbb{E}\Big[\|v(t,\mathbf{X}_{t};\mathbf{y})\|_{2}^{2}-2\langle v(t,\mathbf{X}_{t};\mathbf{y}),\,\mathbf{U}\rangle\Big].$
With the dual architecture, we compute this loss separately for geometry and color and sum them:
$\mathcal{L}_{\mathrm{FM}}=\mathcal{L}_{\mathrm{FM}}^{\text{xy}}+\mathcal{L}_{\mathrm{FM}}^{\text{rgb}},$
where $\mathcal{L}_{\mathrm{FM}}^{\text{xy}}$ is evaluated over all $N$ points, and $\mathcal{L}_{\mathrm{FM}}^{\text{rgb}}$ is evaluated on the single RGB vector per cloud.
Geodesic-length loss (default endpoint loss).
In addition, we employ a shortest-path regularizer using the default endpoint loss (as in our configuration):
$\mathcal{L}_{\mathrm{geo}}=\mathbb{E}\big[\|\mathbf{X}_{1}-\mathbf{X}_{0}\|_{2}^{2}\big].$
In the dual setting, this decomposes as
$\mathcal{L}_{\mathrm{geo}}=\mathcal{L}_{\mathrm{geo}}^{\text{xy}}+\mathcal{L}_{\mathrm{geo}}^{\text{rgb}},\qquad\mathcal{L}_{\mathrm{geo}}^{\text{xy}}=\mathbb{E}\big[\|\mathbf{X}_{1,\text{xy}}-\mathbf{X}_{0,\text{xy}}\|_{2}^{2}\big],\quad\mathcal{L}_{\mathrm{geo}}^{\text{rgb}}=\mathbb{E}\big[\|\mathbf{c}_{1}-\mathbf{c}_{0}\|_{2}^{2}\big].$
Overall training objective.
We optimize the combined objective
$\mathcal{L}=\mathcal{L}_{\mathrm{FM}}+\lambda\,\mathcal{L}_{\mathrm{geo}},$
with $\lambda$ controlling the strength of the geodesic-length regularization.
B.5 GMM implementation
Dataset Model Generated vs. Target Source vs. Target
MMD $\downarrow$ W1 $\downarrow$ W2 $\downarrow$ MMD $\downarrow$ W1 $\downarrow$ W2 $\downarrow$
Letter H (Validation) CFM 0.02021 $\pm$ 0.00115 0.7044 $\pm$ 0.0222 0.8626 $\pm$ 0.0229 0.09731 $\pm$ 0.00101 1.2101 $\pm$ 0.0056 1.3342 $\pm$ 0.0066
SP-FM 0.01387 $\pm$ 0.00092 0.5939 $\pm$ 0.0190 0.7372 $\pm$ 0.0198 0.02773 $\pm$ 0.00042 0.7631 $\pm$ 0.0053 0.9048 $\pm$ 0.0047
Letter W (Test) CFM 0.01778 $\pm$ 0.00101 0.6118 $\pm$ 0.0187 0.7219 $\pm$ 0.0203 0.06088 $\pm$ 0.00049 0.8972 $\pm$ 0.0062 1.0151 $\pm$ 0.0071
SP-FM 0.01352 $\pm$ 0.00143 0.5429 $\pm$ 0.0196 0.6441 $\pm$ 0.0167 0.01872 $\pm$ 0.00062 0.6140 $\pm$ 0.0091 0.7173 $\pm$ 0.0099
Letter Y (Test) CFM 0.02648 $\pm$ 0.00230 0.6669 $\pm$ 0.0295 0.7779 $\pm$ 0.0390 0.06020 $\pm$ 0.00128 1.0614 $\pm$ 0.0048 1.2407 $\pm$ 0.0055
SP-FM 0.01735 $\pm$ 0.00167 0.5406 $\pm$ 0.0259 0.6358 $\pm$ 0.0295 0.04134 $\pm$ 0.00221 0.9525 $\pm$ 0.0399 1.1096 $\pm$ 0.0448